OpenStudy (anonymous):

In right triangle ABC, tanA=0.4. In right triangle DEF, tanD=0.4 △ABC must be similar to △DEF. a.) true b.) false

OpenStudy (anonymous):

If both triangles have the same tan, does that mean that they are similar?

OpenStudy (ankit042):

are the two triangle right angled?

OpenStudy (ankit042):

I am not sure if this is 100% correct. tan for an angle in a triangle is generally defined for right angled triangles so we can use SAS property to prove triangles are similar.

OpenStudy (mathstudent55):

In \( \color{red}{\text{right}} \) triangle ABC, tanA=0.4. In \( \color{red}{\text{right}} \) triangle DEF, tanD=0.4 △ABC must be similar to △DEF. a.) true b.) false

OpenStudy (mathstudent55):



OpenStudy (anonymous):

Okie dokie! So this would make it true, correct?

OpenStudy (anonymous):

Thank you for the very nice visual! Excellent help!

OpenStudy (ankit042):

@mathstudent55 thanks for help! i forgot to look the right angel in the question

OpenStudy (mathstudent55):

From the figure, we have \( \dfrac{BC}{AC} = \dfrac{EF}{DF} \) Sqaring both sides: \( \dfrac{(BC)^2}{(AC)^2} = \dfrac{(EF)^2}{(DF)^2} \) From a property of proportions, we have: \( \dfrac{(BC)^2 + (AC)^2}{(AC)^2} = \dfrac{(EF)^2 + (DF)^2}{(DF)^2} \) From the right triangles, we have: \( (AC)^2 + (BC)^2 = (AB)^2 \) \( (DF)^2 + (EF)^2 = (DE)^2 \) By substitution, we get: \( \dfrac{(AB)^2}{(AC)^2} = \dfrac{(DE)^2}{(DF)^2} \) By a property of proportions, we get: \( \dfrac{(AC)^2}{(DF)^2} = \dfrac{(AB)^2}{(DE)^2} \) Take square roots: \( \dfrac{AC}{DF} = \dfrac{AB}{DE} \) \( \dfrac{BC}{AC} = \dfrac{EF}{DF} \) \( \dfrac{BC}{EF} = \dfrac{AC}{DF} \) \( \dfrac{BC}{EF} = \dfrac{AC}{DF} = \dfrac{AB}{DE}\) Since the lengths of all sides of the one triangle are proportional to the lengths of all corresponding sides of the other triangle, the triangles are similar.

OpenStudy (mathstudent55):

Yes, it's true.

OpenStudy (anonymous):

Thank you for teaching me how to thoroughly figure this out!! You're awesome! =)

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (mathstudent55):

I think @ankit042 got it right way above with using SAS Similarity.

OpenStudy (anonymous):

Yes, good job @ankit042!

OpenStudy (ankit042):

Thanks and you are welcome. @catsarecool1995 you know how to use SAS or need explination?

OpenStudy (anonymous):

Yes thank you very much!!♥